# Multiple shells in supernova 2023ixf support the jittering jets explosion mechanism (JJEM)

### v1

@@ Abstract
Examining the photospheric time evolution of the core-collapse supernova (CCSN) SN 2023ixf from the literature, we identify three evolutionary time periods with constant expansion velocities, which we attribute to three ejecta shells. We find that several CCSN remnants have morphologies with two or more complete or partial shells, compatible with the presence of two or three photospheric shells during the photospheric phase of the explosion. Studies have attributed these CCSN remnants to the jittering-jet explosion mechanism (JJEM), which involves two or three energetic pairs of jets participating in the explosion. We, therefore, conclude that the structure of the photospheric shells of SN 2023ixf supports its explosion by the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
@@ 1 Introduction
A major question about core-collapse supernovae (CCSNe) is the mechanism of their explosion. There are two intensively studied competing theoretical explosion mechanisms of core-collapse supernovae (CCSNe), the delayed neutrino mechanism and the jittering jets explosion mechanism (JJEM).
@@
The research on the delayed neutrino explosion mechanism focuses on simulating the revival of the stalled shock around the newly-born neutron star (NS) at $r\simeq 150~\rm{km}$ by neutrino heating, finding the conditions for an explosion rather than a ‘failed supernova’, and comparing simulated outcomes with observations (e.g., Bamba et al. 2025; Boccioli et al. 2025; Boccioli & Roberti 2025; Eggenberger Andersen et al. 2025; Fang et al. 2025; Huang et al. 2025; Imasheva et al. 2025; Laplace et al. 2025b; Maltsev et al. 2025; Maunder et al. 2024; Mori et al. 2025; Müller et al. 2025; Nakamura et al. 2025; Sykes & Müller 2025; Janka 2025a; Orlando et al. 2025; Paradiso & Coughlin 2025; Tsuna et al. 2025; Vink et al. 2025; Wang & Burrows 2025; Willcox et al. 2025; Mukazhanov 2025; Raffelt et al. 2025; Vartanyan et al. 2025). The magnetorotational explosion mechanism operates in very rare cases of rapid pre-collapse core rotation, leading to the launch of one pair of jets along a fixed axis (e.g., Shibata et al. 2025). At best, it can account for only a small fraction of CCSNe, and studies of the magnetorotational explosion mechanism attribute most CCSNe to the neutrino-driven mechanism; therefore, we group the magnetorotational mechanism with the neutrino-driven mechanism. [^1] See Janka 2025b for a recent talk on the neutrino driven mechanism: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.46.mp4; also https://www.youtube.com/watch?v=nRfDPPSmnzI&t=100s
@@
The research of the JJEM has focused since 2024 on finding supporting evidence for jets in CCSN remnants (CCSNRs; Bear et al. 2025; Bear & Soker 2025; Shishkin et al. 2025; Soker 2025a, b, c, d; Soker & Akashi 2025; Soker & Shishkin 2025a, b, for papers since 2025) [^2] See Soker 2025e For a talk on the JJEM: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.47.mp4. The morphologies of many CCSNRs, in particular point-symmetric CCSNRs (those with two or more symmetry axes), strongly support the JJEM (see the summary of results in, e.g., Soker 2024a, 2025f). Other studies of the JJEM since 2025 are three-dimensional simulations of jet shaping (Braudo et al., 2025) and pre-collapse angular momentum fluctuations in the inner core (Wang et al., 2025).
@@
The explosion process of CCSNe provides only a few observables to distinguish between the two mechanisms. A minority of CCSNe that have explosion energies of $E_{\rm ex}\gtrsim 2\times 10^{51}~\rm{erg}$ support the JJEM because the neutrino-driven mechanism struggles to reach these explosion energies (for recent reviews see Soker 2024a, 2025f). This is the case, for example, in many superluminous stripped-envelope supernovae (e.g., Kumar 2025). However, the two explosion mechanisms have similar predictions to most other observables of the explosion process itself (before the remnant is spatially resolved; Soker 2024a, 2025f). In this study, we find that SN 2023ixf has a property that is more compatible with the JJEM.
@@
We use the photosphere radius evolution $R_{\rm ph}(t)$ of SN 2023ixf that Zimmerman et al. (2024) calculated (Section 2) to claim that jets exploded SN 2023ixf. In Section 3, we present CCSNRs that have jet-shaped morphologies that are compatible with two or more photospheric shells (depending on the line of sight) as we deduced in Section 2. In this study, shells are not necessarily spherical; they can be caps of protrusions termed ears or partial shells, such as the dense boundary of a jet-inflated bubble. We summarize this short study in Section 4.
@@ 2 Identifying photospheric shells in SN 2023ixf
SN 2023ixf is a well-studied nearby CCSN (e.g., Soraisam et al. 2023; Bostroem et al. 2024, 2025; Van Dyk et al. 2024a; DerKacy et al. 2025; Jacobson-Galán 2025; Jacobson-Galán et al. 2025; Medler et al. 2025; Park et al. 2025; Vinko et al. 2025; Zheng et al. 2025). It attracted attention from the beginning because of its proximity and dense, compact circumstellar material (e.g., Berger et al. 2023; Bostroem et al. 2023; Grefenstette et al. 2023; Jacobson-Galán et al. 2023; Smith et al. 2023; Teja et al. 2023; Laplace et al. 2025a). There are uncertainties regarding the progenitor and explosion energy. Singh et al. (2024) model the lightcurve with hydrodynamical calculations and estimated a progenitor radius of $R_{\ast}\simeq 470R_{\odot}$ and an explosion energy of $E_{\rm exp}\simeq 2\times 10^{51}~\rm{erg}$. However, they commented that these values are not unique due to modeling degeneracies. Van Dyk et al. 2024b estimated the progenitor stellar radius to be in the range of $\simeq 921-2012R_{\odot}$, with a median of $R_{\ast}=1389R_{\odot}=9.7\times 10^{13}~\rm{cm}$. Hsu et al. (2025) estimated the progenitor radius to be $R_{\ast}\gtrsim 950R_{\odot}$. Moriya & Singh (2024) estimated the explosion energy to be $E_{\rm exp}\simeq(2-3)\times 10^{51}~\rm{erg}$, while Michel et al. (2025) estimated a more typical energy of $E_{\rm exp}\simeq(0.3-1.4)\times 10^{51}~\rm{erg}$.
@@
We examine the photospheric radius $R_{\rm ph}(t)$ that Zimmerman et al. (2024) calculated in their thorough analysis of SN 2023ixf and give in a table: Supplementary Data (Source Data Table 1). Figure 1 presents the photospheric radius that Zimmerman et al. (2024) calculated at early days (stars and pentagons) as in their Figure 2. We use the same time scale as Zimmerman et al. (2024), who took it from Yaron et al. (2023), but note that the explosion time is uncertain. The time $t=0$ corresponds to the first observed brightening due to the explosion. Yaron et al. (2023) estimated the uncertainty in the time to be about one hour. The explosion at the center of the star has started earlier. In Figure 2, we present the photospheric radii that Zimmerman et al. (2024) calculated at later times.
@@
Figure 1: The time evolution of the photospheric radius from Zimmerman et al. (2024) based on photometry (not on spectroscopy). We fit two straight lines through two groups of points: the first four points, green stars, at $t=1.13-1.42~\rm{day}$ that we mark as Shell 1, and the next 11 points, green pentagons, at $t=1.57-6.13~\rm{day}$ that we mark as Shell 2. The red pentagons are excluded from the fitting; they belong to the phase when the photosphere is moving inward in the mass coordinate of the ejecta. We also indicate the velocity of each shell during the fitting period. The line for Shell 3 fits a later time period, as shown in Figure 2. We identify these shells as photospheric shells appearing one after the other. The straight lines for the three shells are $R_{\rm ph1}=1.050t+0.425$, $R_{\rm ph2}=0.808t+0.477$, and $R_{\rm ph3}=0.445t+0.723$, where the time is in days and the radius in units of $10^{14}~\rm{cm}$. All lines have a coefficient of determination (R-Squared) of $R^{2}=0.98$.
@@
Figure 2: Similar to Figure 1, but including later times, and different scales on the axes. As in 1, red pentagons and pluses are excluded from fittings. We fit a third straight line to the points in the time period of $t=42.19-70.63~\rm{day}$ (green pluses); this is Shell 3.
@@
Zimmerman et al. (2024) argued for a constant photosphere until $t\simeq 2.75~\rm{day}$. We instead find that there are two photospheric shells during that time. We fit two straight lines to the points in Figure 1. Shell 1 that we fit through the first four point (green stars) corresponds to a velocity of $v_{\rm S1}=12,152\pm 4256~\rm{km}~\rm{s}^{-1}$, and Shell 2 that we fit through the green pentagons, to $v_{\rm S2}=9,350\pm 598~\rm{km}~\rm{s}^{-1}$. In Figure 2, we extend the time, and find a time period, represented by the green pluses, where the photosphere increases linearly, this is Shell 3 with a velocity of $v_{\rm S3}=5,152\pm 1214~\rm{km}~\rm{s}^{-1}$. Shell 1 expands, and the photosphere moves outward with it. As it becomes optically thin (transparent), the photosphere expands more slowly as it recedes in the ejecta’s mass coordinate. In the transition from Shell 1 to Shell 2, the photosphere also decreases in radius, suggesting that Shell 1 carries little mass. Then the slower, denser Shell 2 catches up with the photosphere and becomes the photosphere, with linear growth. Later, Shell 2 becomes optically thin, and the photosphere expansion slows down again as it recedes in the ejecta. At about $t\simeq 45~\rm{day}$, Shell 3 catches up and makes the photosphere. The photosphere expands linearly until Shell 3 becomes optically thin at $t\simeq 70~\rm{day}$.
We make the following comments.
(1) From Figure 2, one might propose that there are two shells, where Shell 1 and Shell 2 are part of the same shell. Shell 3 is the same. We find that a straight line fits the first 22 points from $t=1.13~\rm{day}$ to $t=17.81~\rm{day}$. Namely, Shell 1, Shell 2, and the following 7 points form one photospheric shell. The velocity of such a massive early shell is $v_{\rm SE}=7,676\pm 130~\rm{km}~\rm{s}^{-1}$. We consider this less likely (but cannot completely reject this possibility), because (a) the four early points are generally above this line while some of the later points are below it, (b) this line misses some points by more that the error bars, and (c) as we present in Figure 1, Shell 1 and Shell 2 are distinguished.
@@
(3) Tylenda (2005) found three shells for the luminous red nova V838 Mon. In the case of V838 Mon, the photosphere decreases in radius between the shells. In our analysis, we find a decrease between Shell 1 and Shell 2. This decrease is covered by two observational points (transition from green stars to green pentagons in Figure 1). In the transition from Shell 2 to Shell 3, the photosphere expansion slows down, but its radius does not decrease; the transition occurs over a relatively long time, $t\simeq 6~\rm{day}$ to $t\simeq 45~\rm{day}$.
(4) At late times, a free shell of ejecta expands at a constant velocity (if it does not collide with a massive CSM), as its kinetic energy is much larger than its thermal energy. This is not true very close to the origin, when the ejecta is hot and thermal energy is convected to kinetic energy (adiabatic cooling). Since the stellar radius is $\approx 10^{14}~\rm{cm}$, the velocity of the shells close to the center, $r\lesssim 2\times 10^{14}$, is not constant. A collision with the dense, compact CSM might slow the first shells, while the conversion of thermal to kinetic energy accelerates them near the center. The study of the shell’s behavior close to the star is a subject for a separate study.

### v2

@@ Abstract
Examining the photospheric time evolution of the core-collapse supernova (CCSN) SN 2023ixf from the literature, we identify two (possibly three) evolutionary time periods with constant expansion velocities, which we attribute to two (or three) ejecta shells. We find that several CCSN remnants have morphologies with two or more complete or partial shells, compatible with the presence of two (or three) photospheric shells during the photospheric phase of the explosion. Studies have attributed these CCSN remnants to the jittering-jet explosion mechanism (JJEM), which involves two or three energetic pairs of jets participating in the explosion. We, therefore, conclude that the structure of the photospheric shells of SN 2023ixf supports its explosion by the JJEM. This study adds to the accumulating evidence that the JJEM is the primary explosion mechanism of CCSNE.
@@ 1 Introduction
A major question about core-collapse supernovae (CCSNe) is the mechanism of their explosion. There are two intensively studied competing theoretical explosion mechanisms of CCSNe, the delayed neutrino mechanism and the jittering jets explosion mechanism (JJEM).
@@
The research on the delayed neutrino explosion mechanism focuses on simulating the revival of the stalled shock around the newly-born neutron star (NS) at $r\simeq 150~\rm{km}$ by neutrino heating, finding the conditions for an explosion rather than a ‘failed supernova’, and comparing simulated outcomes with observations (e.g., Bambaetal2025CasA; Bocciolietal2025; BoccioliRoberti2025; EggenbergerAndersenetal2025; FangQetal2025; Huangetal2025; Imashevaetal2025; Laplaceetal2025; Maltsevetal2025; Maunderetal2025; Morietal2025; Mulleretal2025; Nakamuraetal2025; SykesMuller2025; Orlandoetal20251987A; ParadisoCoughlin2025; PowellMuller2025; Tsunaetal2025; Vinketal2025; WangBurrows2025; Willcoxetal2025; Mukazhanov2025; Raffeltetal2025; Vartanyanetal2025; Janka2025 for a recent review). The magnetorotational explosion mechanism operates in very rare cases of rapid pre-collapse core rotation, leading to the launch of one pair of jets along a fixed axis (e.g., Shibataetal2025). At best, it can account for only a small fraction of CCSNe, and studies of the magnetorotational explosion mechanism attribute most CCSNe to the neutrino-driven mechanism; therefore, we group the magnetorotational mechanism with the neutrino-driven mechanism. [^1] See Janka2025Padova for a recent talk on the neutrino-driven mechanism: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.46.mp4; also https://www.youtube.com/watch?v=nRfDPPSmnzI&t=100s
@@
The research of the JJEM has focused since 2024 on finding supporting evidence for jets in CCSN remnants (CCSNRs; Bearetal2025Puppis; BearSoker2025; Shishkinetal2025S147; Soker2025G0901; Soker2025N132D; Soker2025RCW89; Soker2025Dust; SokerAkashi2025; SokerShishkin2025Vela; SokerShishkin2025W49B, for papers since 2025) [^2] See Soker2025Padova For a talk on the JJEM: https://www.memsait.it/videomemorie/volume-2-2025/VIDEOMEM_2_2025.47.mp4. The morphologies of many CCSNRs, in particular point-symmetric CCSNRs (those with two or more symmetry axes), strongly support the JJEM (see the summary of results and reviews in Soker2024UnivReview; Soker2025Learning). Other studies of the JJEM since 2025 are three-dimensional simulations of jet shaping (Braudoetal2025) and pre-collapse angular momentum fluctuations in the inner core (WangShishkinSoker2025).
@@
The explosion process of CCSNe provides only a few observables to distinguish between the two mechanisms. A minority of CCSNe that have explosion energies of $E_{\rm ex}\gtrsim 2\times 10^{51}~\rm{erg}$ support the JJEM because the neutrino-driven mechanism struggles to reach these explosion energies (for recent reviews see Soker2024UnivReview; Soker2025Learning). This is the case, for example, in many superluminous stripped-envelope supernovae (e.g., Kumar2025). However, the two explosion mechanisms have similar predictions to most other observables of the explosion process itself (before the remnant is spatially resolved; Soker2024UnivReview; Soker2025Learning). In this study, we find that SN 2023ixf has a property that is more compatible with the JJEM.
@@
We use the photosphere radius evolution $R_{\rm ph}(t)$ of SN 2023ixf that Zimmermanetal2024 calculated (Section 2) to claim that jets exploded SN 2023ixf. In Section 3, we present CCSNRs that have jet-shaped morphologies that are compatible with two or more photospheric shells (depending on the line of sight) as we deduced in Section 2. In this study, shells are not necessarily spherical; they can be caps of protrusions termed ears or partial shells, such as the dense boundary of a jet-inflated bubble. We summarize this short study in Section 4.
@@ 2 Identifying photospheric shells in SN 2023ixf
SN 2023ixf is a well-studied nearby CCSN (e.g., Hiramatsuetal2023; Soraisametal2023; Bostroemetal2024; Bostroemetal2025; VanDyketal2024B; DerKacyetal2025; JacobsonGalan2025; JacobsonGalanetal2025; Medleretal2025; Parketal2025; Vinkoetal2025; Zhengetal2025). It attracted attention from the beginning because of its proximity and dense, compact circumstellar material (e.g., Bergeretal2023; Bostroemetal2023; Grefenstetteetal2023; JacobsonGalanetal2023; SmithNetal2023; Tejaetal2023; Laplaceetal2025B). There are uncertainties regarding the progenitor and explosion energy. Singhetal2024 model the lightcurve with hydrodynamical calculations and estimated a progenitor radius of $R_{\ast}\simeq 470R_{\odot}$ and an explosion energy of $E_{\rm exp}\simeq 2\times 10^{51}~\rm{erg}$. However, they commented that these values are not unique due to modeling degeneracies. VanDyketal2024 estimated the progenitor stellar radius to be in the range of $\simeq 921-2012R_{\odot}$, with a median of $R_{\ast}=1389R_{\odot}=9.7\times 10^{13}~\rm{cm}$. Hsuetal2025 estimated the progenitor radius to be $R_{\ast}\gtrsim 950R_{\odot}$. MoriyaSingh2024 estimated the explosion energy to be $E_{\rm exp}\simeq(2-3)\times 10^{51}~\rm{erg}$, while Micheletal2025 estimated a more typical energy of $E_{\rm exp}\simeq(0.3-1.4)\times 10^{51}~\rm{erg}$.
@@
We examine the photospheric radius $R_{\rm ph}(t)$ that Zimmermanetal2024 calculated in their thorough analysis of SN 2023ixf and give in a table: Supplementary Data (Source Data Table 1). Figure 1 presents the photospheric radius that Zimmermanetal2024 calculated at early days (stars and pentagons) as in their Figure 2. We use the same time scale as Zimmermanetal2024, who took it from Yaronetal2023, but note that the explosion time is uncertain. The time $t=0$ corresponds to the first observed brightening due to the explosion. Yaronetal2023 estimated the uncertainty in the time to be about one hour. The explosion at the center of the star has started earlier. In Figure 2, we present the photospheric radii that Zimmermanetal2024 calculated at later times.
@@
Figure 1: The time evolution of the photospheric radius from Zimmermanetal2024 based on photometry (not on spectroscopy). We fit two straight lines through two groups of points: the first four points, green stars, at $t=1.13-1.42~\rm{day}$ that we mark as Shell 1, and the next 11 points, green pentagons, at $t=1.57-6.13~\rm{day}$ that we mark as Shell 2. The red pentagons are excluded from the fitting; they belong to the phase when the photosphere is moving inward in the mass coordinate of the ejecta. We also indicate the velocity of each shell during the fitting period. The line for Shell 3 fits a later time period, as shown in Figure 2. We identify these shells as photospheric shells appearing one after the other. The straight lines for the three shells are $R_{\rm ph1}=1.050t+0.425$, $R_{\rm ph2}=0.808t+0.477$, and $R_{\rm ph3}=0.445t+0.723$, where the time is in days and the radius in units of $10^{14}~\rm{cm}$. All lines have a coefficient of determination (R-Squared) of $R^{2}=0.98$. The velocities and their uncertainties are from fitting the ejecta photosphere radii without including the radii of the shells at $t=0$. Including the radii at $t=0$ changes the velocities little but substantially reduces the uncertainties, indicating that the three shells are distinct, by these photospheric radii (caveat: see text on the possibility that Shells 1 and 2 are a single shell).
@@
Figure 2: Similar to Figure 1, but including later times, and different scales on the axes. As in Figure 1, red pentagons and pluses are excluded from fittings. We fit a third straight line to the points in the time period of $t=42.19-70.63~\rm{day}$ (green pluses); this is Shell 3. The points from day 7 to day 41 (red pentagons) are the photospheric radius in the ejecta inward to Shell 2 and outside Shell 3. During that time, the photosphere moves inward in the mass coordinate of the ejecta. Therefore, each point corresponds to a deeper mass layer than the previous point; the points do not represent a single layer. For that, we cannot fit a straight line through these points. Around day 45, the front of Shell 3 becomes the photosphere. From there to about day 71, a straight line corresponds to the front of Shell 3. After day 71, Shell 3 became transparent, and the photosphere moved inward in the mass coordinate of the ejecta (red pluses).
@@
Zimmermanetal2024 argued for a constant photosphere until $t\simeq 2.75~\rm{day}$. We instead find that there are two photospheric shells during that time. We fit two straight lines to the points in Figure 1. Shell 1 that we fit through the first four point (green stars) corresponds to a velocity of $v_{\rm S1}=12,152\pm 4256~\rm{km}~\rm{s}^{-1}$, and Shell 2 that we fit through the green pentagons, to $v_{\rm S2}=9,350\pm 598~\rm{km}~\rm{s}^{-1}$; these lines are the fit through the points of the ejecta photosphere only, not including the photosphere (stellar radius) at $t=0$. In Figure 2, we extend the time, and find a time period, represented by the green pluses, where the photosphere increases linearly, this is Shell 3 with a velocity of $v_{\rm S3}=5,152\pm 1214~\rm{km}~\rm{s}^{-1}$ as fitted through the points of photospheric radii but not through the center. The three fitted lines through the ejecta photospheric radii $R_{\rm ph}(t)$ cross the expected stellar radius at $t=0$, as should be for expanding ballistic (more or less) shells.
Shell 1 expands, and the photosphere moves outward with it. As it starts to become transparent (optically thin), the photosphere expands more slowly as it recedes in the ejecta’s mass coordinate. In the transition from Shell 1 to Shell 2, the photosphere also decreases in radius, suggesting that Shell 1 carries little mass. Then the slower, denser Shell 2 catches up with the photosphere and becomes the photosphere, with linear growth. Later, Shell 2 becomes optically thin, and the photosphere expansion slows down again as it recedes in the ejecta. At about $t\simeq 45~\rm{day}$, Shell 3 catches up and makes the photosphere. The photosphere expands linearly until Shell 3 becomes optically thin at $t\simeq 71~\rm{day}$.
We make the following comments: (1) When we draw the first two shells just by the photospheric points, 4 points for Shell 1 and 11 points for Shell 2, the velocities have large uncertainties $v_{\rm S1}=12,152\pm 4256~\rm{km}~\rm{s}^{-1}$ and $v_{\rm S2}=9,350\pm 598~\rm{km}~\rm{s}^{-1}$. However, this fit is intended to show that the straight lines pass through the center of the explosion. To derive the shell velocity with the correct uncertainties, we must include the explosion’s center. We do not know exactly the stellar radius or the time the shells broke from the photosphere. We set the same initial radius for the three shells at $t=0$, rather than a different initial radius for each shell, although it is clear they did not all pass through that radius at the same time. For a shells’ radius of $R_{6}(0)=650R_{\odot}\pm 200R_{\odot}$ at $t=0$ the velocities change little, but the uncertainties are much smaller now: $v_{\rm S1r6}=11,911\pm 1,231~\rm{km}~\rm{s}^{-1}$, $v_{\rm S2r6}=9,391\pm 477~\rm{km}~\rm{s}^{-1}$, and $v_{\rm S3r6}=5213\pm 193~\rm{km}~\rm{s}^{-1}$; for $R_{7}(0)=750R_{\odot}\pm 250R_{\odot}$ the velocities are $v_{\rm S1r7}=11,381\pm 1483~\rm{km}~\rm{s}^{-1}$, $v_{\rm S2r7}=9,280\pm 507~\rm{km}~\rm{s}^{-1}$, and $v_{\rm S3r7}=5,197\pm 195~\rm{km}~\rm{s}^{-1}$; Subscripts $r6$ and $r7$ denote the two shells’ radii at $t=0$ that we use, respectively. For larger initial shells’ radii than $\simeq 750R_{\odot}$ (at $t=0$), the velocities of Shells 1 and 2 depart more from the velocities of the points alone (the velocities given in Figures 1 and 2), and so we regard these shells’ radii at $t=0$ as less likely. Shells 1 and 2 are about one sigma apart, and Shells 2 and 3 are about six sigma apart. These velocities, along with their uncertainties, reinforce the hypothesis that the system comprises of at least two shells (1+2 as one shell, and Shell 3), and with some uncertainty from three shells.
(2) From Figure 2, one might propose that there are two shells, where Shell 1 and Shell 2 are part of the same shell. Shell 3 is the same as before. We find that a straight line fits the first 22 points from $t=1.13~\rm{day}$ to $t=17.81~\rm{day}$. Namely, Shell 1, Shell 2, and the following 7 points form one photospheric shell. The velocity of such a massive early shell is $v_{\rm SE}=7,676\pm 130~\rm{km}~\rm{s}^{-1}$. This velocity is very similar to the $8,000~\rm{km}~\rm{s}^{-1}$ early-time ejecta velocity that Zimmermanetal2024 deduce. The velocity similarity supports a massive shell. In addition, the first four points that make Shell 1 do not have UV observations, and therefore, the fitting of the black-body temperature and radius is less certain. Given the photospheric radii of Shell 1 as calculated by Zimmermanetal2024, we consider the interpretation of Shell 1 and 2 as separate shells more likely. However, if the temperatures of these four points are underestimated and their photospheric radii are smaller, causing them to lie along the line of Shell 2, then Shell 1 and 2 would merge into a single Shell. In that case, the interpretation of Zimmermanetal2024 of a constant photosphere until $t\simeq 2.75~\rm{day}$ does not hold. Overall, we caution that Shell 1 and Shell 2 may not be separate shells, but rather a single Shell. Even under this interpretation, in which Shells 1 and 2 are one massive shell, Shell 3 is separate, and the shells model exists with two shells.
@@
(4) Tylenda2005 found three shells for the luminous red nova V838 Mon. In the case of V838 Mon, the photosphere decreases in radius between the shells. In our analysis, we find a decrease between Shell 1 and Shell 2. This decrease is covered by two observational points (transition from green stars to green pentagons in Figure 1). In the transition from Shell 2 to Shell 3, the photosphere expansion slows down, but its radius does not decrease; the transition occurs over a relatively long time, $t\simeq 6~\rm{day}$ to $t\simeq 45~\rm{day}$. Namely, the transition from Shell 2 to Shell 3 is gradual, and the photosphere moves inward only slightly in mass coordinate before Shell 3 catches up. This gradual and small deepening of the photosphere in mass coordinate explains the small change in the slope of $R_{\rm ph}(t)$ between Shell 2 and Shell 3 (Figure 2), and the continuous slope in SN 2023ixf luminosity in that phase.
(5) At late times, a free shell of ejecta expands at a constant velocity (if it does not collide with a massive circumstellar matter), as its kinetic energy is much larger than its thermal energy. This is not true very close to the origin, when the ejecta is hot and thermal energy is convected to kinetic energy (adiabatic cooling). Since the stellar radius is $\approx 10^{14}~\rm{cm}$, the velocity of the shells close to the center, $r\lesssim 2\times 10^{14}$, is not constant. A collision with the dense, compact circumstellar matter might slow the first shells, while the conversion of thermal to kinetic energy accelerates them near the center. The study of the shell’s behavior close to the star is a subject for a separate study.